OpenStudy (anonymous):

A wall with no windows is 11 feet high and 20 feet long. A large roll of wall paper costs $25 and will cover 60 square feet of a wall. A small roll of wall paper cost $6 and will cover 10 square feet of wall. What is the least cost for enough wall paper to cover the wall?

OpenStudy (anonymous):

@hartnn @UditKulka

OpenStudy (anonymous):

Total sq ft= 11*20 =220 sqft for large roll..cost per sqft= 25/60 for small roll it is 6/10 To cover up, lets consider larger roll required for x sqft wall and small roll require y sq ft wall lest also consider Z is the cost hence Z= (25/60)*x + (6/10)*y where x+y=220 z= [25*x+36*y]/60 for least cost we need to minimize z.. to do that we need larger value of x than the value of y (as co-efficient of x less than that of y) hence x must be greater than 110. now 60*z= 25*x + 36(220-x)= 36*220 -11*x z= 132 -11*x/60 z= 132- 11*180/60 z=99 hence 99

OpenStudy (anonymous):

OMGGGGG :o Thanks alot :)